OpenStudy (anonymous):

I really really need some help tonight. PLEASE. I am to plot the points between and beyond the x-intercept and the vertical asymptotes and then type the y-coordinate for each of the following points. The formula is f(x)= x+2/x^2+x-6 The following numbers are the points given in a graph beside x in cells: -5 -1, 0, 1,5 Can anyone help or give me a link to a calculator that will accomplish this? None of the ones I have tried can.

OpenStudy (anonymous):

seems as if everyone has fallen asleep

OpenStudy (anonymous):

yeap it's really easy

OpenStudy (anonymous):

How so? :-D

OpenStudy (anonymous):

y=u/v-->y'=(u'v-v'u)/v2

OpenStudy (anonymous):

then for figure out the critical points solve the equation of y'=0

OpenStudy (anonymous):

it can show the y'=0 is going to be u/v=u'/v'

OpenStudy (anonymous):

or x+2/x2+x-6=1/2x+1 then 2x2+5x+2=x2+x-6 or x2+4x+8=0 there is no any real answer

OpenStudy (anonymous):

the curve has two vertical saymptotic x=-3 and x=2

OpenStudy (anonymous):

And also it has one horizentally asymptotic y=0

OpenStudy (anonymous):

Do you know why?

OpenStudy (anonymous):

Not really.

OpenStudy (anonymous):

because on x=-3 and x=2 y is going to infinite

OpenStudy (anonymous):

and if x approach to infinite y is going to zero

OpenStudy (anonymous):

but domain id bounded between[-5 5]

OpenStudy (anonymous):

you can draw a table and plot the curve

OpenStudy (anonymous):

Still not getting it but thank you for explaining. :-)